Use the quick formula to find the approximated duration from a 10 basis point change in yield, note you will need to find thePlease solve and show work!
Answer #1

Solution) Coupon = 3%

YTM = 2.5%

Maturity (years) = 4

Par value = $100

Bond quoted at Price per par (V0) = $101.8920

Face Value = $1000

Change in Yield, \Delta y = 10 basis point = 10/10000 = 0.001 = 0.1%

First, increase the yield on the bond (YTM) by 10 basis points from 2.5% to 2.6%. The new price (V+), quoted on par, is computed as:

Year Cash Flows
t=1 $                    3
t=2 $                    3
t=3 $                    3
t=4 $               103
YTM 2.60%
Bond Current Price   =NPV(YTM, Cash flows)
$         101.5012

Thus, V+ = $101.50

Second, decrease the yield on the bond (YTM) by 10 basis points, from 2.5% to 2.4%. The new price (V-), quoted on par, is computed as:

Year Cash Flows
t=1 $                    3
t=2 $                    3
t=3 $                    3
t=4 $               103
YTM 2.40%
Bond Current Price =NPV(YTM, Cash flows)
$         102.2626

Thus, V- = $102.26

Approximate modified duration is calculated using the below formula:

Modified Duration = (V_-V) 2(V)(Ay)

where, \Delta y = change in yield = 0.1% = 0.001

Thus, Approximate modified duration = (102.2626 - 101.5012)/2*101.8920*0.001 = 0.7614/0.203784 = 3.7363